OpenStudy (anonymous):

The area of a triange with base,b and height h, is given by A=1/2bh If the area of a certain triangeis 65 square inches,what is the base?

OpenStudy (anonymous):

You at least need the variable h to solve this, I believe.

OpenStudy (anonymous):

thats what i was thinking when the instructor assign this problem seems like it was missing a number.

OpenStudy (anonymous):

Must be a mistake on their part then, because I cant seem to find any other way to solve it, sorry. Keep this bumped, maybe someone else knows.

OpenStudy (anonymous):

65 = (1/2)bh b = (65 x 2)/h b = 130/h @ragerun is correct you need the height

OpenStudy (anonymous):

a different height gives us a different answer for base

OpenStudy (anonymous):

Yeah, and it's possible your instructor wants the answer with the variable for some reason. In that case use the answer @eigenschmeigen got which is 130h for the base.